one fifth of the female students and one tenth of the male students eat in cafeteria.if there are a total of 800 female students and 600 male students what fraction of the students do not eat in the cafeteria?
Answer by fractalier(6550)

One-fifth of the 800 female students is (1/5)(800) = 160 students.
One-tenth of the 600 males students is (1/10)(600) = 60 students.
Therefore 220 students eat in the cafeteria.
That means (800+600) - 220 = 1400 - 220 = 1180 students do not.
1180/1400 = 84.3% do not eat in the cafeteria.